OpenStudy (moongazer):

In a circle diameter of 8in. find the arc length subtented on the circumference by a central angle of 90degrees. Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (sriram):

u surely know to make the question attractive!

OpenStudy (sriram):

circumference=2*pi*r=2*3.14*8 inches=50.265 angle at center =90 degrees hence if u imagine 1/4 th of the circumference is involved hence arc length=50.265/4=12.566

OpenStudy (anonymous):

90 is a quarter of the circle so the arc is a quarter of the circumference.